40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1889     Entry time: Thu Jan 7 17:09:16 2021
Author: Paco & Anchal 
Type: DailyProgress 
Category: Optics 
Subject: Mode matching OPO 

Fresh attempt at mode matching. For this,

  1. Installed the oven, plugged it to the controller and went to the nominal temperature setpoint (40 C) to match the expected path length inside the NL crystal
  2. Placed the output coupler (roc = 15 mm) and roughly align so that the retroreflection is overlapped with the input beam.
  3. Set up a PD (Newfocus 2001) and scope, operating the laser at relatively low power (current ~ 840 mA), and optimize the FI rejected power.
  4. The output coupler is mounted on a three axis mirror mount (Polaris, hoping to get low drift) such that we have some knobs to tune the mode matching initially.

After a couple of iterations moving the mirror X,Y and then scanning all knobs (X,Y, and XY) to effectively translate along Z, the optimized FI rejection is ~(2.15 mW /2.95 mW) 75% of the input beam power. Looking closely at the backreflection from the output coupler, I can clearly see multiple scattered spots, which could definitely account for the defficiency. The most likely culprit is the crystal itself, which is mounted between brass and glass surfaces with no respect for anti-reflection measures. The waist is small enough that no clipping should be happening, so it looks like the NL crystal placement may have to be revisited. Other than that, this procedure should be fine.

ELOG V3.1.3-